### v1

@@ 8 Understanding and Interpreting SVERL
SVERL provides a lens through which we can understand how individual feature values contribute to an agent’s behaviour, performance, and value estimation. These insights go beyond what the policy or expected return reveals alone. While many of the conclusions drawn in our examples may seem intuitive in hindsight, they were not obvious when observing only the agent’s actions or value function. Even in well-understood domains, SVERL revealed how specific features shaped decision-making and outcomes, offering a more detailed understanding of the reinforcement learning problem.
@@
In this section, we reflect on the insights provided by SVERL, considering how the three explanation types compare and contrast to form a richer picture of agent behaviour. We also discuss how these explanations should be interpreted in practice, including common pitfalls and the risk of misunderstanding.
8.1 Comparing Explanations
@@
Each explanation in the proposed framework captures a distinct aspect of reinforcement learning. Explanations of behaviour reveal how features influence the agent’s action choices. Explanations of performance reveal how features influence expected return. Explanations of value estimation reveal how features influence the estimation of the expected return. Understanding how these perspectives agree—and when they diverge—is important for forming a general understanding of agent-environment interactions.
The final state of Mastermind shows an example of when explanations of behaviour and performance align. The same features that drive the agent’s decision to guess AB also determine the outcome: they reveal the correct code and lead to the highest possible return. This agreement may not be surprising, since performance is determined by behaviour. However, SVERL also demonstrates that this alignment between behaviour and performance is not guaranteed. In the initial state of Mastermind, the empty bottom row contributes to the first guess but has no impact on performance, since all initial guesses yield the same return. This highlights a subtle but important point: a feature can shape the agent’s behaviour without affecting the return that behaviour produces.
The initial state of Mastermind also shows an example of explanations of behaviour and value estimation mirroring each other. In the initial state of Mastermind, the empty bottom row contributes positively to selecting the first guess and negatively to the return estimate. These attributions have the same relative magnitudes but opposite signs, reflecting that the feature increases the likelihood of an action while identifying a low-return state. In contrast, the same features contribute positively to both explanations in the final state: they identify the correct action and a high-return state. This pattern arises because both explanations are defined over the steady-state distribution—they depend on the information a feature provides about the underlying state, which informs both the agent’s behaviour and the return expected from that state.
Both performance and value estimation explain expected return, but differ in how features influence it. Performance explanations attribute responsibility via the agent’s behaviour—what the agent does changes when features are unknown. Value estimation explanations, in contrast, reflect how features inform return estimates without altering behaviour. As shown in the initial state of Mastermind, the bottom row contributes nothing to performance but negatively to value estimation: although it doesn’t change the return received, it signals that the agent is in a low-return state.
@@ 8.2 Interpreting SVERL
As with any feature attribution method, interpreting SVERL requires care. While Shapley values reveal how features influence behaviour, performance, or return estimates, they do not necessarily explain why. Human users naturally form beliefs or hypotheses based on these values, but it is easy to overinterpret or misattribute causality. Making these values useful for explanation often depends on insights from psychology and cognitive science, which help bridge the gap between attribution and understanding. SVERL provides a foundation for such interpretation, but further analysis is needed to assess whether the conclusions are well-founded.
Figure 4: Feature contributions to the probability of opening each available square in the two states of Minesweeper shown in Figure 3. The colour of each square shows its Shapley value (scaled between $-1$ and $1$) for contributing to the probability of opening that same square.
Consider the example in Minesweeper. In Figure 3, SVERL assigns negative performance attributions to unopened squares that contain mines in the second state. One might conclude that observing these features lowers performance because it increases the likelihood that the agent opens them. This is a reasonable hypothesis, but it must be validated. Figure 4 shows the contribution of each square to the probability of the agent opening that square. In both states, the unopened squares contribute positively to the probability of being opened. This supports the hypothesis, though it was not a guaranteed conclusion from the original explanation alone.
Misunderstandings can also arise when explaining contributions by focusing too narrowly on individual marginal terms within the Shapley summation. An example of this is the misinterpretation of the $x$ and $y$ contributions in the Taxi domain by Beechey et al. (2023), where these terms were accidentally swapped (corrected in Section B.4). While the qualitative discussion was not entirely incorrect, it misrepresented the full contributions by considering only a subset of feature interactions, overlooking that Shapley values are means over all possible feature coalitions. This highlights the importance of interpreting Shapley-based explanations in light of the full combinatorial structure they represent.
@@ 9 Related Work
We now situate the proposed framework within the broader landscape of explainable reinforcement learning. We first review prior work that has applied Shapley values to analyse reinforcement learning agents, before discussing other lines of research in explainability.
9.1 Previous Applications of Shapley Values in Reinforcement Learning
While Shapley values have previously been applied to explain reinforcement learning agents, this work is the first to study them mathematically in this setting, focusing on the validity and meaning of the resulting explanations.
Several prior works analyse the value function using SHAP (Lundberg and Lee, 2017), a Python package that uses marginal characteristic functions to compute Shapley values for models (Zhang et al., 2020; Schreiber et al., 2021; Zhang et al., 2021). Applied to the value function, this setup corresponds to a marginal version of SVERL’s value estimation explanation, which assumes feature independence. However, as shown throughout this paper, value estimation concerns the prediction of return, not how features influence the agent’s behaviour or the return the agent actually receives. Interpreting such explanations as relating to behaviour risks misunderstanding the agent’s interaction with its environment.
Figure 5: Tic-Tac-Toe. Left: feature contributions to the agent’s return estimate (value estimation). Right: contributions to the agent’s performance. The agent (X) plays against a minimax opponent (O) on a $3\times 3$ grid, aiming to place three marks in a row. Players take turns and have an equal probability of starting. Each state is defined by nine features, corresponding to the contents of the grid squares.
This distinction is illustrated in Figure 5, where an optimal agent plays as X against a minimax (Polak, 1989) opponent in Tic-Tac-Toe. In the state shown, the agent must block the opponent’s winning move. The performance explanation correctly attributes responsibility to the opponent’s marked squares. Value estimation, however, assigns zero contribution to all features. This is because the optimal agent consistently forces a draw against the minimax opponent. As a result, the expected return remains fixed at the value corresponding to a draw, independent of the state and, consequently, the values of its features. For an optimal policy, there is no statistical association between feature values and return estimates. While value estimation can identify informative features for prediction, it does not explain behaviour or performance.
@@ 9.1 Previous Applications of Shapley Values in Reinforcement Learning
Finally, Wäldchen et al. (2022) compute Shapley values by learning characteristic functions through training the policy and the value function with masked features. This approach restricts the number of masked features to stabilise learning, undermining the Shapley principle of evaluating all possible feature subsets. This is an important limitation because a feature’s importance may only emerge when enough others are removed. This limitation prevents the method from satisfying the Shapley axioms and producing principled attributions, unlike SVERL, which is axiomatically sound.
@@ 9.2 Other Approaches to Explainable Reinforcement Learning
Many existing approaches to explaining reinforcement learning agents aim to identify important input features or clarify decision behaviour. A widely studied class of methods is based on saliency, where explanations are derived by perturbing inputs and observing changes in outputs (Greydanus et al., 2018; Puri et al., 2020) or by computing input gradients (Wang et al., 2016). While often useful in practice, especially in visual domains, these methods lack theoretical guarantees and can produce inconsistent explanations for different models that yield identical outputs. By contrast, SVERL is grounded in the Shapley framework, ensuring consistency and offering a principled, model-agnostic foundation for attributing importance.
@@
Other approaches analyse MDP structure and policies more broadly. Some leverage knowledge of transition dynamics to simulate counterfactuals (Rupprecht et al., 2019), identify causal influences (Madumal et al., 2020), or estimate task completion probabilities (Cruz et al., 2019, 2023). Others rely on assumptions about reward decomposability to trace how different objectives contribute to value estimates (Juozapaitis et al., 2019) or identify critical states where incorrect decisions would lead to poor performance (Amir et al., 2018; Huang et al., 2018). These methods can offer useful insights but typically depend on explicit assumptions about environment models or reward structure, or are limited in scope, highlighting states without explaining why actions are chosen. By contrast, SVERL provides general-purpose, instance-level behaviour explanations without making such assumptions.
@@ 10.1 Explaining Agent-Environment Interactions More Broadly
We have focused on explaining agent behaviour at specific states or for specific actions. Understanding behaviour broadly often involves reasoning about trends across many interactions. While one could present the Shapley values for many individual states and actions, doing so imposes a high cognitive load and hinders comprehension (Covert et al., 2021). Here, we explore how SVERL could be extended to summarise explanations across states or actions to provide broader insights.
A natural idea is to aggregate SVERL’s explanations for individual states or actions by taking their expectation under some distribution $\mathcal{D}$. For example, given a behaviour explanation $\phi_{i}(\tilde{\pi}_{s}^{a})$ for action $a$ at state $s$, we compute:
@@
as a general framework for summarising feature contributions. While we illustrate this approach for explanations of behaviour, the same approach applies to performance and value estimation. The choice of distribution $\mathcal{D}$ determines the kind of insight obtained. It could reflect, for example, the full range of the agent’s experience, focus on particularly salient states, or highlight trends conditional on specific feature values.
@@
which represents the average contribution of feature $i$ to the probability of action $a$ across the agent’s experience, weighting each feature contribution by the likelihood it’s observed. However, perhaps counterintuitively, this quantity is always zero.
@@
This result follows from the efficiency axiom: Shapley values explain deviations from a baseline, which is defined as the expected action probability over $p^{\pi}(s)$. On average, these deviations must sum to zero over $p^{\pi}(s)$. A similar argument applies to value estimation, as stated below.
@@
For any policy $\pi$, the expected contribution of any feature to the return estimate under the steady-state distribution is zero:
@@
In contrast, performance explanations can yield non-zero expectations:
@@
which represents the average contribution of feature $i$ to the probabilities of actions in state $s$, weighting each feature contribution by the likelihood that the action is chosen. One could interpret this concept as capturing the general influence of a feature value on behaviour across actions at a given state. By the linearity of Shapley values, this quantity is itself a Shapley value for a new coalitional game with characteristic function defined as the expected action probability:
@@ 10.2 Approximating SVERL
Computing SVERL’s explanations is computationally demanding. Each characteristic function involves an expectation over the entire state space, and each Shapley value is defined as a sum over all $2^{|{\cal F}|}$ subsets of features. These requirements quickly become intractable as the number of features or states grows. As a result, practical applications of SVERL in large domains require approximation techniques.
Approximation techniques have been widely studied in the supervised learning setting, and it is only natural to expect that many of these methods could be extended to SVERL. Here, we adapt one such method, the Monte Carlo sampling techniques in Section 2.5, to approximate SVERL’s explanations of behaviour, performance, and value estimation. We present this theoretical outline to demonstrate that SVERL can be approximated in practice, leaving its full treatment and discussion of alternative methods to future work.
We begin with explanations of behaviour and value estimation. We propose approximating their characteristic functions, defined in Equations 8 and 11, using Monte Carlo sampling:
@@
Performance explanations require a different treatment. The characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$, defined in Equation 10, is the expected return from state $s$ under a policy that follows the partially observed policy $\tilde{\pi}_{s}^{a}(\mathcal{C})$ in $s$, and the fully observed policy $\pi$ elsewhere.
Because this characteristic is an expected return, it can be estimated using any standard reinforcement learning method. However, it depends on the policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$, which must also be approximated. We propose incorporating the sampling-based approximation of $\tilde{\pi}_{s}^{a}(\mathcal{C})$ from Equation 13 directly into the policy definition used within the performance characteristic’s expectation:
@@
This defines a new policy $\hat{\pi}_{\mathcal{C}}(a_{t}\mid s_{t})$ that, when visiting $s$, selects actions based on a sampled state $s^{\prime}$ consistent with the observed features: $s_{\mathcal{C}}=s^{\prime}_{\mathcal{C}}$, and follows the fully observed policy $\pi$ elsewhere. By rewriting the performance characteristic in this equivalent form, we embed the approximation of the policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$ into the return estimation itself, allowing $\tilde{v}^{\pi}_{s}(\mathcal{C})$ to be computed directly using standard reinforcement learning methods.
Finally, with $\tilde{v}^{\pi}_{s}(\mathcal{C})$ approximated in this way, the Shapley values for performance can be computed using the usual permutation-based sum:
@@
All three explanations—behaviour, performance, and value estimation—require access to the conditional steady-state distribution $p^{\pi}(S\mid S_{\mathcal{C}}=s_{\mathcal{C}})$. This distribution must be learned in large domains, presenting an important direction for future work. One promising starting point is the parametric approach introduced by Frye et al. (2020a), which uses variational inference to learn conditional data distributions from a dataset. A natural adaptation to the reinforcement learning setting would be to use this method to learn the conditional steady-state distribution from a replay buffer of agent experience.
@@
Here we have presented an outline for approximating SVERL by adapting Monte Carlo techniques from supervised learning to the reinforcement learning setting, without investigating empirical performance or efficiency. We encourage a full treatment of these questions in subsequent research. Exploring other approximation methods developed in supervised learning may also prove fruitful. For example, Frye et al. (2020a) propose learning parametric models of the characteristic function directly, and Jethani et al. (2021) amortises Shapley value approximations across inputs by training parametric models to predict them. These approaches are promising future avenues to enable more scalable or efficient approximations of SVERL.
10.3 Using Value Estimation to Inform Behaviour

### v2

@@ 8 Understanding and Interpreting SVERL
SVERL provides a lens through which we can understand how individual feature values contribute to an agent’s behaviour, outcomes, and predictions. These insights go beyond what the policy or expected return reveal alone. While many of the insights from our examples may seem intuitive in hindsight, they were not obvious when observing only the agent’s actions or value function. Even in well-understood domains, SVERL revealed how specific features shaped decision-making, outcomes, and predictions, offering a more detailed understanding of reinforcement learning.
@@
In this section, we reflect on the insights provided by SVERL. We first examine how the three explanation types offer complementary perspectives that together form a richer picture of agent-environment interactions. We then discuss interpreting these explanations in practice, including common pitfalls and the risk of misunderstanding.
8.1 SVERL’s Complementary Insights
@@
Each explanation in the proposed framework captures a distinct aspect of reinforcement learning. Explanations of behaviour reveal how features influence the agent’s action choices. Explanations of outcomes reveal how features influence the consequences of the agent’s behaviour, such as expected return. Explanations of predictions reveal how features influence the estimation of outcomes, such as estimated expected return. While each perspective offers insight, considering them together provides a more complete understanding of agent-environment interactions.
Consider the initial Mastermind state in Figure 3. Here, the behaviour and prediction explanations assign importance to the same features with equal magnitudes but opposite signs, reflecting that the information these features convey about the underlying state simultaneously increases the likelihood of the agent’s action while decreasing its estimate of expected return. In contrast, the outcome explanation attributes no importance to these features, illustrating a subtle but important point: features can influence behaviour and be used to predict outcomes without affecting the outcomes themselves. By comparison, the same features influence all three explanations in the final state. They identify the optimal action, which leads to the highest return—aligning behaviour, outcomes, and predictions in a single coherent interpretation.
These relationships illustrate how SVERL’s explanations, though distinct in focus, can meaningfully interact. They highlight how features play layered roles in shaping decisions, outcomes, and estimates.
@@ 8.2 Interpreting SVERL
As with any attribution method, interpreting SVERL requires care. Shapley values quantify the contribution of each feature to a particular explanatory target, such as an action probability, expected return, or value estimate. They do not reveal causal mechanisms or the internal reasoning behind the agent’s choices. Human users often construct explanations based on these attributions, but doing so can lead to overinterpretation or misplaced causal inference. Beyond their mathematical foundations, for Shapley-based explanations to be useful in practice, insights from cognitive science and psychology can help bridge the gap between statistical association and human understanding. SVERL provides a principled basis for attribution, but further analysis can help ensure that its outputs support meaningful and useful explanations in particular domains.
Figure 5: Feature contributions to the probability of opening each available square in the two states of Minesweeper shown in Figure 4. The colour of each square shows its Shapley value (scaled between 1-1- 1 and 111) for contributing to the probability of opening that same square.
Consider the example in Minesweeper. In Figure 4, SVERL assigns negative outcome attributions to unopened squares that contain mines in the second state. One might conclude that observing these features lowers expected return because it increases the likelihood that the agent opens them. This is a reasonable hypothesis, but it must be validated. Figure 5 shows the contribution of each square to the probability of the agent opening that square. In both states, the unopened squares contribute positively to the probability of being opened. This supports the hypothesis, though it was not a guaranteed conclusion from the original explanation alone.
Misunderstandings can also arise when explaining contributions by focusing too narrowly on individual marginal terms within the Shapley summation. An example of this is the misinterpretation of the $x$ and $y$ contributions in the Taxi domain by Beechey et al. (2023), where these terms were accidentally swapped (corrected in Appendix B). While the qualitative discussion was not entirely incorrect, it misrepresented the full contributions by considering only a subset of feature interactions, overlooking that Shapley values are means over all possible feature coalitions. This highlights the importance of interpreting Shapley-based explanations in light of the full combinatorial structure they represent.
@@ 9 Related Work
We now situate the proposed framework within the broader landscape of explainable reinforcement learning. We first review prior work that has applied Shapley values to analyse reinforcement learning agents, then discuss other lines of research in explainability.
9.1 Previous Use of Shapley Values in Reinforcement Learning
While Shapley values have previously been used to explain reinforcement learning agents, this work is the first to study them mathematically in this setting, focusing on the validity and meaning of the resulting explanations.
Several prior works analyse the value function using SHAP (Lundberg and Lee, 2017), a Python package that uses marginal characteristic functions to compute Shapley values for models (Zhang et al., 2020; Schreiber et al., 2021; Zhang et al., 2021). Applied to the value function, this setup corresponds to a marginal version of SVERL’s prediction explanation, which assumes feature independence. However, as shown throughout this paper, prediction concerns the estimation of expected return, not how features influence the agent’s behaviour or the expected return the agent actually receives. Interpreting such explanations as relating to behaviour risks misunderstanding the agent’s interaction with its environment.
This distinction is illustrated in the Tic-Tac-Toe example of Figure 2 in Section 6.4. In that example, the outcome explanation intuitively attributes responsibility to the opponent’s marked squares. Prediction, however, assigns zero contribution to all features because all states lead to a draw with expected return zero. There is no statistical association between feature values and return estimates; predictions can identify informative features for estimating return, but it does not explain how those features shape behaviour or outcomes.
@@ 9.1 Previous Use of Shapley Values in Reinforcement Learning
Finally, Wäldchen et al. (2022) compute Shapley values by learning characteristic functions through training the policy and the value function with masked features. This approach restricts the number of masked features to stabilise learning, undermining the Shapley principle of evaluating all possible feature subsets. This is an important limitation because a feature’s importance may only emerge when enough other features are removed. This limitation prevents the method from satisfying the Shapley axioms and producing principled attributions, unlike SVERL, which is axiomatically sound.
@@ 9.2 Other Approaches to Explainable Reinforcement Learning
Feature importance methods are a common approach to explaining reinforcement learning agents, aiming to quantify how state features influence agent behaviour. These methods are particularly popular in visual domains. In addition to Shapley-based approaches, common alternatives include perturbing inputs and observing changes in outputs (Greydanus et al., 2018; Puri et al., 2020) or computing input gradients (Wang et al., 2016). However, these methods lack the theoretical guarantees provided by Shapley values. For example, some can produce inconsistent attributions across models with identical outputs (Sundararajan et al., 2017). In contrast, SVERL is grounded in the Shapley framework, ensuring consistency and offering a principled, model-agnostic foundation for attributing importance.
@@
Other approaches analyse MDP structure and policies more broadly. Some leverage knowledge of transition dynamics to simulate counterfactuals (Rupprecht et al., 2019), identify causal influences (Madumal et al., 2020), or estimate task completion probabilities (Cruz et al., 2019, 2023). Others rely on assumptions about reward decomposability to trace how different objectives contribute to value estimates (Juozapaitis et al., 2019) or identify critical states where incorrect decisions would lead to poor performance (Amir et al., 2018; Huang et al., 2018). These methods can offer useful insights but depend on explicit assumptions about environment models or reward structure, or are limited in scope, highlighting states without explaining why actions are chosen. By contrast, SVERL provides general-purpose, instance-level behaviour explanations without making such assumptions.
@@ 10.1 Explaining Agent-Environment Interactions More Broadly
We have focused on explaining agents at specific states. Understanding behaviour broadly often involves reasoning about trends across many interactions. While one could present the Shapley values for many individual states and actions, doing so imposes a high cognitive load and hinders comprehension (Covert et al., 2021). Here, we explore how SVERL could be extended to summarise explanations across states or actions to provide broader insights.
A natural idea is to aggregate SVERL’s explanations for individual states or actions by taking their expectation under some distribution $\mathcal{D}$. For example, given a Shapley value from a behaviour explanation $\phi_{i}(\tilde{\pi}_{s}^{a})$ for action $a$ at state $s$, we compute:
@@
as a general framework for summarising feature contributions. While we illustrate this approach for explanations of behaviour, the same approach applies to outcomes and predictions. The choice of distribution $\mathcal{D}$ determines the kind of insight obtained. It could reflect, for example, the full range of the agent’s experience, focus on particularly salient states, or highlight trends conditional on specific feature values.
@@
which represents the mean contribution of feature $i$ to the probability of action $a$ across the agent’s experience, weighting each feature contribution by the probability that it is observed. However, perhaps counterintuitively, this quantity is always zero.
@@
This result follows from the efficiency axiom: Shapley values explain deviations from a baseline, which is defined as the expected action probability over $p^{\pi}(s)$. On average, these deviations must sum to zero over $p^{\pi}(s)$. A similar argument applies to predictions, as stated below.
@@
For any policy $\pi$, the expected contribution of any feature to the estimate of expected return under the steady-state distribution is zero:
@@
In contrast, outcome explanations can yield non-zero expectations:
@@
which represents the mean contribution of feature $i$ to the probabilities of actions in state $s$, weighting each feature contribution by the probability that the action is chosen. One could interpret this concept as capturing the general influence of a feature value on behaviour across actions at a given state. By the linearity of Shapley values, this quantity is itself a Shapley value for a new coalitional game with characteristic function defined as the expected action probability:
@@ 10.2 Approximating SVERL
Computing SVERL’s explanations is computationally demanding. Each characteristic function involves an expectation over the entire state space, and each Shapley value is defined as a sum over all $2^{|{\cal F}|}$ subsets of features, where ${\cal F}$ is the full set of features. These requirements quickly become intractable as the number of features or states grows. As a result, practical applications of SVERL in large domains require approximation techniques.
Approximation techniques have been widely studied in the supervised learning setting, and it is only natural to expect that many of these methods could be extended to SVERL. Here, we adapt one such method, the Monte Carlo sampling techniques in Section 2.5, to approximate SVERL’s explanations of behaviour, outcomes, and predictions. We present this theoretical outline to demonstrate that SVERL can be approximated in practice, leaving its full treatment and discussion of alternative methods to future work.
We begin with explanations of behaviour and predictions. We propose approximating their characteristic functions, defined in Equations 8 and 11, using Monte Carlo sampling:
@@
Outcome explanations require a different treatment. The characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$, defined in Equation 10, is the expected return from state $s$ under a policy that follows the partially observed policy $\tilde{\pi}_{s}^{a}(\mathcal{C})$ in $s$, and the fully observed policy $\pi$ elsewhere.
Because this characteristic is an expected return, it can be estimated using any standard reinforcement learning method. However, it depends on the discrete policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$, which must also be approximated. We propose incorporating the sampling-based approximation of $\tilde{\pi}_{s}^{a}(\mathcal{C})$ from Equation 13 directly into the policy definition used within the outcome characteristic’s expectation:
@@
This defines a new policy $\hat{\pi}_{\mathcal{C}}(a_{t}\mid s_{t})$ that, when visiting $s$, selects actions based on a sampled state $s^{\prime}$ consistent with the observed features: $s_{\mathcal{C}}=s^{\prime}_{\mathcal{C}}$, and follows the fully observed policy $\pi$ elsewhere. By rewriting the outcome characteristic in this equivalent form, we embed the approximation of the discrete policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$ into the return estimation itself, allowing $\tilde{v}^{\pi}_{s}(\mathcal{C})$ to be computed directly using standard reinforcement learning methods.
Finally, with $\tilde{v}^{\pi}_{s}(\mathcal{C})$ approximated in this way, the Shapley values for outcomes can be computed using the usual permutation-based sum:
@@
All three explanations—behaviour, outcomes, and predictions—require access to the conditional steady-state distribution $p^{\pi}(S\mid S_{\mathcal{C}}=s_{\mathcal{C}})$. This distribution must be learned in large domains, presenting an important direction for future work. One promising starting point is the parametric approach introduced by Frye et al. (2020a), which uses variational inference to learn conditional data distributions from a dataset. A natural adaptation to the reinforcement learning setting would be to use this method to learn the conditional steady-state distribution from a replay buffer of agent experience.
@@
Here we have presented an outline for approximating SVERL by adapting Monte Carlo techniques from supervised learning to the reinforcement learning setting, without investigating empirical performance or efficiency. We encourage a full treatment of these questions in subsequent research. Exploring other approximation methods developed in supervised learning may also prove fruitful. For example, Frye et al. (2020a) propose learning parametric models of the characteristic function directly, and Jethani et al. (2021) amortises Shapley value approximations across inputs by training parametric models to predict them. These approaches are promising future avenues to enable scalable or efficient approximations of SVERL.
10.3 Using Predictions to Inform Behaviour